Floyd can walk 2 mph faster than his wife, Rachel. It takes Rachel 3 hr longer than Floyd to hike a 12-mi trail through the park. Find their speeds.
Answer by jorel555(1290)

Let r be the rate at which Rachel walks. Then Floyd is r+2. So:
12/(r+2)+3=12/r
12r+3rē+6r=12r+24
3rē+6r-24=0
rē+2r-8=0
(r+4)(r-2)=0
r=2, -4
Throwing out the negative answer, we get Rachel's rate at 2 mph, and Floyd's at 4 mph. ☺☺☺☺